OpenStudy (anonymous):

Determine whether –4 is a solution of –9x – 6 = 42.

OpenStudy (anonymous):

-4 is not a solution because if you plug it into x you would get this: -9(-4)-6=42 36-6=42 30=42 obviously 30 is not equal to 42 so therefore -4 is not equal to x

OpenStudy (anonymous):

that make a sense thank you

OpenStudy (anonymous):

No problem:) Do you need to find what is equal to x you would go like this: -9x-6=42 -9x=48 x=-16/3 or -5.33

OpenStudy (anonymous):

thanks I write down that it help me to work out some of the problem